OpenStudy (anonymous):

16. A painter leans a 20-ft ladder against a building. The base of the ladder is 12 ft from the building. To the nearest foot, how high on the building does the ladder reach

OpenStudy (anonymous):

Use the Pythagorean theorem to find building height. The ladder is c

OpenStudy (anonymous):

a^2 + 12^2=20^2

OpenStudy (anonymous):

23??

OpenStudy (anonymous):

a^2 + 144=400 a^2= 400-144 a^2= 256 a= 16 ft